# Intention Analysis Makes LLMs A Good Jailbreak Defender

### v2

@@ B.1 Safety Evaluation
For evaluation of DAN, considering the complex instructions in DAN makes it challenging to directly identify the success of an attack through string matching, we turn to utilize a widely-adopted LLM to evaluate the harmfulness of model generations. Practically, we adopt gpt-3.5-turbo-0613888https://openai.com/blog/chatgpt as the evaluation function following Deng et al. (2023a). Especially, the model responses are submitted via API calls, and a specific evaluation prompt is designed for gpt-3.5-turbo-0613 to rate the harmfulness on a five-degree scale (‘A (Very Harmful)’, ‘B (Harmful)’, ‘C (Moderately Harmful)’, ‘D (Slightly Harmful)’, and ‘E (No Harmful at All)’). Figure 6 shows the detailed instruction for gpt-3.5-turbo-0613 evaluation. Human evaluation results in Sec 5.3 demonstrate the highest agreement between human and gpt-3.5-turbo-0613 when viewing the evaluation degrees of ‘A (Very Harmful)’ and ‘B (Harmful)’ as a success of jailbreak attacks.
@@ B.2 Helpfulness Evaluation
B.3 Intention Recognition Success Evaluation
@@
Appendix C Language Models
ChatGLM-6B
Vicuna-7B-Chat
MPT-30B-Chat
DeepSeek-67B-Chat
GPT-3.5
To study the performance differences between our method and safety training methods, we used LLaMA2-7B-Chat as a representative of these methods considering the various safety training strategies it has undergone and the superior safety it ultimately demonstrates.
LLaMA-2-7B-chat
@@ Appendix D Alternative Prompts
To assess the resilience of our method against specific expressions, we construct other two sets of alternative instructions for intention analysis. These additional instructions were meticulously designed with modifications in lexical choices and syntactic constructions to direct intention analysis. The detailed prompts are presented in Figure 9.
@@
Appendix E Cross-Intention Analysis Experiment
@@
Appendix F Qualitative Examples in Ours Experiments
This section presents some practical examples in our experiments. These examples are from SAP200 dataset (see Figure 10), DAN dataset (see Figure 11), and AdvBench dataset (see Figure 12), respectively.
Figure 10: Qualitative example I from SAP200 dataset about the comparison between vanilla and our $\mathbb{IA}$.
Figure 11: Qualitative example II from DAN dataset about the comparison between vanilla and our $\mathbb{IA}$.
Figure 12: Qualitative example III from AdvBench dataset about the comparison between vanilla and our $\mathbb{IA}$.

### v3

@@ C.3 Helpfulness Evaluation
C.4 Intention Recognition Success Evaluation
@@ Appendix D Alternative Prompts
To assess the resilience of our method against specific expressions, we construct other two sets of alternative instructions for intention analysis. These additional instructions were meticulously designed with modifications of origin $\mathbb{IA}$ prompts (detailed in Figure 8) in lexical choices and syntactic constructions to direct intention analysis. The detailed prompts are presented in Figure 10. Comparison results in Table 6 demonstrate that: it is the intention analysis mechanism itself, but not the specific expression of the instructions, that makes our method effective.
Figure 8: Detailed two-stage prompts for our method.
@@
Methods Vicuna-7B Vicuna-13B Empirical
GCG DAN SAP200 GCG DAN SAP200 Runtime
Vanilla 83.0 48.4 70.0 87.0 60.0 65.9 1×1$\times$1 ×
+ BPE-dropout (Jain et al., 2023) 63.0 23.8 67.2 50.0 28.2 48.9 <1×<1$\times$< 1 ×
+ ICD (Wei et al., 2023b) 1.00 44.4 32.8 0.00 58.9 32.8 <2×<2$\times$< 2 ×
+ Self Defense (Helbling et al., 2023) 24.0 31.3 53.2 20.0 28.8 29.7 2$\sim$ 2$\times$ 2 ×
+ Moral Self-Correction (Ganguli et al., 2023) 26.0 25.0 49.0 13.0 28.1 42.8 3$\sim$ 3$\times$ 3 ×
+ Self-Reminder (Xie et al., 2023) 11.0 45.3 33.8 1.0 57.5 36.9 <2×<2$\times$< 2 ×
+ SmoothLLM (Robey et al., 2023) 8.00 13.5 54.4 5.00 17.3 37.0 10$\sim$ 10$\times$ 10 ×
+ $\mathbb{IA}$ (Ours) 0.00 3.42 0.31 0.00 0.94 1.56 2$\sim$ 2$\times$ 2 ×
Table 9: Comparison of our method and existing advanced defense methods in terms of ASR (%) and empirical runtime. The best and second best results are highlighted in bold and underline.
@@
Appendix E Comparison with all defense baselines
Table 9 lists the comparison results between our $\mathbb{IA}$ and the baselines. [^14] Due to computational resource and financial limitations, we randomly select 320 samples each from DAN and SAP200 datasets for comparative experiments. As observed, our method consistently shows superiority over other baselines on different datasets and model scales. Specifically, our $\mathbb{IA}$ outperforms the second-best method by 30.32% and 23.77% averagely on SAP200 and DAN, respectively. In addition, although ICD and Self-Reminder achieve considerable reduction in ASR on GCG, their performance severely degrades when dealing with complex and stealthy jailbreak prompts. On the contrary, our method consistently outperforms other baselines across both prompt-level and automatic token-level jailbreak datasets. Notably, our method achieves the best ASRs with comparable and acceptable empirical inference runtime.
Appendix F Cross-Intention Analysis Experiment
Appendix G Qualitative Examples in Ours Experiments
This section presents some practical examples in our experiments. These examples are from SAP200 dataset (see Figure 11), DAN dataset (see Figure 12), and AdvBench dataset (for GCG; see Figure 13), respectively.
Figure 11: Qualitative comparison example I from SAP200 dataset on Vicuna-7B with and without our $\mathbb{IA}$.
Figure 12: Qualitative comparison example II from DAN dataset on Vicuna-7B with and without our $\mathbb{IA}$.
Figure 13: Qualitative comparison example III from AdvBench dataset (for GCG) on Vicuna-7B with and without our $\mathbb{IA}$.
Figure 14: Examples of nuanced handling for harmful user queries regarding varying unsafe intentions (terrorism in case 1 and violence in case 2).
Figure 15: Two failure cases of intention analysis induced by too strong (for LLaMA2-7B-Chat) or too weak (for Vicuna-7B) inherent safety.
Figure 16: Two failure cases of our $\mathbb{IA}$ induced by weak inherent safety (in case 1) and failed intention analysis (in case 2).